Hi
A basket of fruit was shared by A B C. A received 50% of the fruit and an extra 4 pieces. B received 40% of the remainder
plus an extra 10 pieces. C received 38 pieces. How many pieces were in the basket at first. Thanks

Answer by ikleyn(52781)

.
A basket of fruit was shared by A B C. A received 50% of the fruit and an extra 4 pieces. B received 40% of the remainder
plus an extra 10 pieces. C received 38 pieces. How many pieces were in the basket at first.
~~~~~~~~~~~~~


Let x = "How many pieces were in the basket at first"  (the unknown under the question).


"A" received  0.5x + 4  pieces.


Remained  x - (0.5x + 4) = 0.5x -4  pieces.


"B" received  0.4*(0.5x-4) + 10  pieces.


Equation for the total is

    (0.5x+4) + 0.4*(0.5x-4) + 10 + 38 = x   pieces.


Simplify and solve

    0.5x + 4 + 0.2x - 1.6 + 10 + 38   = x

    4 -1.6 + 10 + 38                  = x - 0.5x - 0.2x

       50.4                           =      0.3x

        x                             = 50.4/0.3 = 168.


ANSWER.  There were 168 pieces in the basket at first.

Solved.